If g(x) = f(x-2) and f(x) = 4 - (x^2), what is the value of g(0)?
Answer by ikleyn(52795)

.

    g(0) = by the definition = f(0 - 2) = f(-2) = by the definition = 4 - ((-2)^2) = 4 - 4 = 0.    ANSWER

Solved and thoroughly explained.